OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). x + y = 100 0.07x + 0.05y = 5.8

OpenStudy (anonymous):

multiply the second equation by 100 and start with \[x+y=100\\ 7x+5y=580\]

OpenStudy (anonymous):

then you can multiply the first equation by \(-5\) to get \[-5x-5y=-500\\ 7x+5y=580\] add them up and get \[2x=80\] so you can solve for \(x\) in one step

OpenStudy (anonymous):

X=40 than?

OpenStudy (anonymous):

So Y =60

OpenStudy (anonymous):

Thanks! :)